The young modulus of elasticity is defined as

E=f/a÷e/l

obtain the dimension of E where f=force

A=area

e and l =length








1
Expert's answer
2021-11-07T19:25:14-0500

"E=\frac{F}{A}/\frac{e}{l}"


"dim[E]=\frac{N}{m^2}/\frac{m}{m}=\frac{N}{m^2}=Pa"